GNU Rush – a restricted user shell

version 2.2, 2 January 2022

Sergey Poznyakoff

Published by the Free Software Foundation, 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301 USA Copyright \begin{matrix}c \\ ◯ \end{matrix} 2008–2022 Sergey Poznyakoff Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Operation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Configuration File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 5 Default Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6 Usage Tips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7 Test Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8 Option Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 9 The rushwho utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 10 The rushlast utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 11 Accounting Database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 12 How to Report a Bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 A Time and Date Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 B GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . 65 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Configuration File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4.1 Lexical Structure of the Configuration File . . . . . . . . . . . . . . . . . . . . . . 7 4.2 Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.3 The global statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4.3.1 Expansion control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3.2 Debugging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.3.3 The sleep-time statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3.4 Error Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3.5 The regexp statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3.6 The include-security statement . . . . . . . . . . . . . . . . . . . . . . . . 13 4.3.7 Accounting control statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

4.4 Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

4.4.1 The Request. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.4.1.1 Positional variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.4.1.2 Request variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.4.1.3 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.4.1.4 User-defined variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.4.1.5 Variable Expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

4.4.2 Matching Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.4.2.1 Comparisons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.4.2.2 Membership operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.4.2.3 File system tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.4.2.4 Boolean expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

4.4.3 Modifying variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.4.3.1 The set statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.4.3.2 The insert statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.4.3.3 The unset statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.4.3.4 The remopt statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.4.3.5 The delete statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.4.3.6 The map statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

4.4.4 Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.4.5 Transformations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4.6 System Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.4.7 Fall-through . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.4.8 Accounting and Forked Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.4.9 Post-process Notification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.4.10 Exit rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.4.11 Interactive Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.4.12 Localization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

4.4.12.1 Localization Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.4.12.2 Writing Your Localization . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.5 Include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

5 Default Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . 39

6 Usage Tips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

6.1 scp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6.2 rsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 6.3 sftp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 6.4 cvs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.5 svn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.6 git . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 6.7 Notification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7 Test Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.1 Dump Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

8 Option Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

9 The rushwho utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

9.1 Rushwho Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 9.2 Output Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

10 The rushlast utility. . . . . . . . . . . . . . . . . . . . . . . . . . . 55

10.1 Rushlast Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

11 Accounting Database . . . . . . . . . . . . . . . . . . . . . . . . . 57

11.1 The wtmp file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 11.2 The utmp file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

12 How to Report a Bug . . . . . . . . . . . . . . . . . . . . . . . . 59

Appendix A Time and Date Formats . . . . . . . . . . . 61 Appendix B GNU Free

Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . 65

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

1 Introduction

GNU Rush is a Restricted User Shell, designed for sites that provide limited remote access to their resources, such as svn or git repositories, scp, or the like. Using a sophisticated configuration file, GNU Rush gives you complete control over the command lines that users execute, as well as over the usage of system resources, such as virtual memory, CPU time, etc.

2 Operation

GNU Rush is usually installed as a user shell. When a user connects to the server (e.g. by using using SSH protocol), the shell binary, rush, is executed. GNU Rush must be called with exactly two arguments: the -c command line option and a command line to be executed on the host machine^{1}. If wrong arguments are supplied, the shell aborts.

The third argument to rush supplies a command line to be executed. That command line, shell environment for its execution and the password database entry for the user who executes rush are said to form a request.

After startup, rush reads a set of rules from its configuration file. Each rule consists of matching conditions and actions. Conditions decide whether the request matches the rule. They can include regular expression matching against entire command line or particular words thereof, comparisons of user name or group,, etc. If all conditions match the request, actions are executed. Actions can instruct rush to:

• Modify the command line;

• Impose resource limits;

• Set umask;

• Change current working directory;

• Modify the execution environment;

• Run command in a special root directory (‘chroot’).

Finally, after all actions have been completed successfully, rush runs the requested command. Notice, that by that time the resulting command line is not necessarily the same as the original one supplied to rush with the -c option.

A special kind of rules, called fall-through ones, is provided. Fall-through rules differ from other rules in that they do not execute the command. After all actions in a fall-through rule have been executed, GNU Rush continues to search for another matching rule in its configuration and applies it, if found. Fall-through rules are useful to set default values for subsequent rules.

1 Starting from version 1.6, it is possible to use GNU Rush for interactive shell sessions.

See Section 4.4.11 [Interactive], page 33, for more information about it.

3 Quick Start

To give you the feel of GNU Rush possibilities, let’s consider the following configuration file rule:

    rush 2.0
    rule sftp
      # Matching condition
      match $uid >= 100 && $command ~ "^.*/sftp-server"
      # Actions:
      set [0] = "bin/sftp-server"
      umask 002
      chroot "~"
      chdir "/"

The first clause defines the version of the syntax this configuration uses. Each configuration must begin with this statement.

Next clause, rule, defines a new rule. Its argument serves as a rule tag and is used for diagnostic messages and in accounting.

Lines beginning with ‘#’ are comments, they are intended for a human reader and are ignored by rush.

The match statement, defines condition that must be met for this rule to become active. In this example it requests that the UID of the requesting user be greater than or equal to 100, and the command line begin with ‘/sftp-server’, optionally preceded by arbitrary directory components.

Subsequent clauses define actions associated with this rule.

The set clause contains instructions on how to modify the first argument of the command line. Argument indices start at 0, so ‘[0]’ refers to the command name. The expression in our example instructs GNU Rush to replace it with ‘bin/sftp-server’.

The umask clause sets the file creation mask.

The chroot clause instructs GNU Rush to chroot to the user home directory before executing the command.

Finally, the chdir statement sets the directory to change to after installing the chroot.

4 Configuration File

The configuration file rush.rc is located in /usr/local/etc by default.^{1}.

The configuration file is read and parsed right after start up. Any errors occurred in parsing are reported using syslog facility ‘authpriv’ and priority ‘notice’. When run in ‘test’ mode, all diagnostics is displayed on the standard error output. See Chapter 7 [Test Mode], page 47, for a detailed description of ways to debug and test your configurations.

Before parsing, rush checks the ownership and permissions of the configuration file for potential security breaches. The configuration file is considered unsafe if any of the following conditions are met:

1. It is not owned by root.

2. It is group writable.

3. It is world writable.

4. It resides in a group writable directory.

5. It resides in a world writable directory.

6. It is a symbolic link to a file residing in a group or world writable directory. If the file is considered unsafe, rush rejects it and aborts execution. Any of these tests can be disabled using the --security-check option

(see [–security-check], page 49).

As of version 2.2, rush supports two distinct configuration file formats. The legacy configuration format is the one used in rush versions up to

1.9. It is still supported to facilitate transition of existing installations to the new syntax. Its support will eventually be removed in future versions, so the users are encouraged to switch to the new syntax as soon as possible. The legacy syntax is described in detail in http://www.gnu.org.ua/

software/rush/legacy.

This manual describes new configuration file format.

4.1 Lexical Structure of the Configuration File

Configuration file consists of tokens separated by arbitrary amount of whitespace characters: horizontal spaces and tabs. Except when enclosed in double quotes or preceded by a dollar sign, the ‘#’ character introduces an inline comment: the character itself and any material that follows it up to the end of the physical line is ignored. Comments are treated as newlines.

The following classes of tokens are recognized.

Newlines A newline character (ASCII 10) terminates a statement. If newline is immediately preceded by a backslash, both characters are removed and the following line is treated as a continuation of the current line. This allows for splitting exceedingly long statements over several physical lines.

1 The exact location of the configuration file is defined when configuring the package.

See the file INSTALL in the GNU Rush source directory for more information

Identifiers Identifiers begin with a letter and consist of letters, digits, underscores and dashes. They serve as keywords and variable names.

Decimal numbers A sequence of decimal digits, optionally preceded by a minus or plus sign.

Unquoted strings An unquoted string is any contiguous sequence of any characters, except newlines, whitespace and the following special characters: ‘\’, ‘"’, ‘!’, ‘=’, ‘<’, ‘>’, ‘(’, ‘)’, ‘{’, ‘}’, ‘[’, ‘]’, ‘$’, ‘%’, ‘&’, ‘|’, ‘~’, ‘#’.

Quoted strings A quoted string is a sequence of characters enclosed in doublequotes. Quoted strings are subject to backslash interpretation, backreference interpretation and variable expansion. During backslash interpretation, the escape sequences are recognized and replaced as per table below:

Sequence  Replaced with
\a  Audible bell character (ASCII 7)
\b  Backspace character (ASCII 8)
\f  Form-feed character (ASCII 12)
\n  Newline character (ASCII 10)

\r

Carriage return character (ASCII 13)

\t

Horizontal tabulation character

(ASCII 9)

\v

Vertical tabulation character

(ASCII 11)

\\  Single backslash character
\"  Double-quote
\%  Percent character

Table 4.1: Backslash escapes

A backslash immediately followed by newline character is removed. A backslash followed by any other character except as listed above is retained along with the character. During backreference interpretation, references to parenthesized groups in regular expression are replaced with the actual content of the corresponding group in the most recently matched string. A reference is ‘%{n}’ where n is a decimal number. If n is one digit, curly braces can be omitted: ‘%n’. If the ‘%’ character results from previous backslash interpretation, no backreference interpretation occurs.

Strings used in the left-hand side of a comparison expression are subject to variable expansion. This is discussed in detail in Section 4.4.1.5 [Variable expansion], page 16.

Variable references Variable references consist of a ‘$’ sign, followed by the positional argument number or variable name, optionally enclosed in curly braces. Positional arguments greater than 9 must be enclosed in curly braces. The variable name must follow the rules for valid identifiers: it must begin with a letter and consist of letters, digits and underscores. Variable name in curly braces can be followed by ‘-’, ‘=’, ‘?’, or ‘+’, optionally preceded by ‘:’ as summarized in the table below:

Reference  Meaning
${var:-word}  Use Default Values
${var:=word}  Assign Default Values
${var:?word}  Display Error if Null or Unset
${var:+word}  Use Alternate Value

Table 4.2: Variable reference

Where word stands for any valid token as described in this section. See Section 4.4.1.5 [Variable expansion], page 16, for a detailed discussion of these forms and their meaning.

Comparison and boolean operators These are:

‘&&’  Boolean and
‘||’  Boolean or
‘!’  Boolean negation
‘==’  Equality (string or numeric)
‘!=’  Inequality (string or numeric)
‘<’  Less than
‘<=’  Less than or equal to
‘>’  Greater than
‘>=’  Greater than or equal to
‘~’  Regexp matching
‘!~’  Negated regexp matching
‘in’  Membership in set of strings
‘group’  Membership in UNIX group
‘=’  Assignment
‘=~’  Regular expression substitution

Table 4.3: Operators

See Section 4.4.2 [Matching Conditions], page 18, for a detailed discussion.

4.2 Syntax

The ‘rush’ configuration consists of statements.

A statement consists of a keyword and optional arguments, separated by any amount of whitespace. Each statement occupies one line in the configuration file and is terminated by a newline character. Extremely long statements may be split across several physical lines by ending each line except the last with a backslash followed by a newline.

Statements may be separated by any amount of empty lines or comments. The first statement in a configuration file indicates the syntax version. It

has the following form:

rush 2.0

This statement is mandatory. In its absence, the file will be treated as a legacy configuration file^{2}. To avoid confusion, a notice message to that effect will be printed.

Statements that follow form logical groups. Each group begins with a rule or global statement.

The global statement introduces global settings. It affects all statements that follow it.

The rule statement introduces a single rush rule, that defines how to process a particular command.

2 For the discussion of the legacy syntax, please refer to http://www.gnu.org.ua/

software/rush/legacy.

These statements are described in the sections that follow.

4.3 The global statement

The global statement defines global settings. The syntax is:

      global
        stmt1
        stmt2
        ...

where dots represent any number of statements. The following subsections discuss the statements that can be used within a global block.

4.3.1 Expansion control

The following statement controls the behavior of rush when an undefined variable is expanded (see Section 4.4.1.5 [Variable expansion], page 16).

[global] expand-undefined bool

If bool is ‘true’, expand undefined variables to empty value. If it is ‘false’ (the default), issue an error and abort. The following values can be used as synonyms for ‘true’: ‘yes’, ‘on’, ‘t’, ‘1’. The following values can be used as synonyms for ‘false’: ‘no’, ‘off’, ‘nil’, ‘0’.

See [handling of undefined variables], page 17, for a detailed discussion of how rush processes undefined variables and for the recommended techniques of handling them.

4.3.2 Debugging

The debug global statement sets the debugging level – an integer value that controls the verbosity of rush:

[global] debug num

Set debugging level to num.

The greater num is, the more verbose is the logging. The debugging information is reported via syslog at facility ‘authpriv’, priority ‘debug’. As of version 2.2, the following debugging levels are supported:

1 A minimum debugging level, and the only one whose messages are logged using the priority ‘notice’. At this level, rush only logs requests and rules selected to handle them. For example:

              rush[16821]: Serving request "/usr/libexec/sftp-server"
              for sergiusz by rule sftp-savane

2 List all actions executed when serving requests.

3 When parsing a legacy configuration file, verbosely describe parsing process.

More debugging levels may be implemented in future.

4.3.3 The sleep-time statement

[global] sleep-time num

Set the time to sleep before exiting on error, in seconds. This statement is intended as a measure against brute-force attacks. Default sleep time is 5 seconds.

4.3.4 Error Messages

[global] message class text

Define a textual message which is returned to the remote party if an error of the given class occurs.

Valid values for class are:

usage-error This error is reported when rush has been invoked improperly. The default text is:

    You are not permitted to execute this command.

nologin-error Define a textual message which is returned to the remote user if there is no such user name in the password database. Default is:

    You do not have interactive login access to this machine.

config-error Define a textual message which is returned to the remote party if the rush configuration file contains errors. Default is:

    Local configuration error occurred.

system-error Define a textual message which is returned to the remote party if a system error occurs. Default message is:

    A system error occurred while attempting to execute command.

4.3.5 The regexp statement

The regexp statement configures the flavor of regular expressions for use by subsequent match, set, and insert statements.

[global] regexp flags ...

Configure the type of regular expressions.

Each flag is a word specifying some regular expression feature. It can be preceded by ‘+’ to enable this feature (this is the default), or by ‘-’ to disable it. Valid flags are:

‘extended’

Use POSIX Extended Regular Expression syntax when interpreting regex. This is the default.

‘basic’ Use basic regular expressions. Equivalent to ‘-extended’.

‘icase’
‘ignore-case’

Do not differentiate case. Subsequent regex matches will be case insensitive.

For example, the following statement enables POSIX extended, case insensitive matching:

 global
   regex +extended +icase

4.3.6 The include-security statement

Additional configuration can be included to the main configuration file using the include statement (see Section 4.5 [Include], page 36). Before inclusion, a number of checks is performed on the file to ensure it is safe to rely on it. These checks are configured using the following statement:

[global] include-security list

Configure the security checks for include files. This statement takes a list of arguments, separated by white space. The following arguments are recognized: all Enable all checks. owner The file is not owned by root. iwgrp groupwritablefile

The file is group writable.

iwoth worldwritablefile

The file is world writable.

dir iwgrp groupwritabledir

The file resides in a group writable directory.

dir iwoth worldwritabledir

The file resides in a world writable directory. link The file is a symbolic link to a file residing in a group or world

writable directory.

Each of the above keywords may be prefixed by ‘no’, which reverses its meaning. The special keyword ‘none’ disables all checks. Each keyword adds or removes a particular test to the existing check list, which is initialized as described in [security checks], page 7. Thus, the following statement results in all checks, except for the file ownership:

    global
      include-security noowner

In the example below, the check list is first cleared by using the none statement, and then a set of checks is added to it:

    global
      include-security none owner iwoth iwgrp

4.3.7 Accounting control statements

The following global statements control file mode and permissions of the accounting database files. For a detailed description of this feature, See Chapter 11 [Accounting Database], page 57.

[global] acct-umask mask

Set umask used when accessing accounting database files. Default value is ‘022’.

[global] acct-dir-mode mode

Set mode bits for the accounting directory. The mode argument is the mode in octal.

[global] acct-file-mode mode

Set mode bits for the wtmp and utmp files.

4.4 Rule

The rule statement configures a GNU rush rule. This is a block statement, which means that all statements located between it and the next rule statement (or end of file, whichever occurs first) modify the definition of that rule.

The syntax of the rule statement is:

[Configuration] rule tag

The tag argument is optional. If it is given, it supplies a tag for the rule, i.e. a (presumably unique) identifier, which is used to label this rule. Rush uses this tag in its diagnostic messages. For rules without explicit tag, Rush supplies a default tag, which is constructed by concatenating ‘#’ character and the ordinal number of rule in the configuration file, in decimal notation. Rule numbering starts from ‘1’.

Each rule group can contain a number of statements that control what kind of requests match that rule and what actions are taken when the rule is matched. Arguments to this statements can refer to command line arguments and other parts of the request.

4.4.1 The Request

User request consists of the user passwd entry, the command line supplied to rush, and environment variables. The request is analyzed and can be eventually modified by rules in rush configuration file. Rules access parts of the request using variables.

There are four classes of variables. All of them share the same namespace and are accessed using the same syntax.

4.4.1.1 Positional variables

Rush performs word splitting using the same rules as sh. Statements in the configuration file refer to command line arguments (words) by their index, using positional variables. A positional variable can have the following forms:

     $n
     ${n}

where n is the variable index. The form with curly braces must be used if n is negative (see below) or greater than 9.

Arguments are numbered from ‘0’. The name of the command is argument ‘$0’. Consider, for example, the following command line:

    /bin/scp -t /upload

Word splitting phase results in three positional variables being defined:

Variable  Value
$0  /bin/scp
$1  -t
$2  /upload

These values can also be referred to using negative indexes. They refer to words in the reverse order, as illustrated in the following table (notice the use of curly braces):

Variable  Value
${-3}  /bin/scp
${-2}  -t
${-1}  /upload

Notice also, that negative indexes are 1-based.

One final note about the ‘$0’ variable. Immediately after word splitting it refers to both the executable program name and the 0th argument that will be passed to that program (argv[0]). Most of the time the two values coincide. However, the rule can modify either value, so that they become different. Whether modified or not, the actual name of the program to be run is kept in the request variable ‘$program’ (see the following section).

4.4.1.2 Request variables

The following variables can be used to refer to various parts of the user request:

Variable  Expansion
$user  User name
$group  Name of the user’s principal group
$uid  UID
$gid  GID
$home  User’s home directory
$gecos  User’s GECOS field
$program  Executable program name
$command  Entire command line
$#  Number of arguments in ‘$command’

4.4.1.3 Environment variables

Environment variables are accessed using the same syntax as the rest of the variables. Rules can modify them using the setenv, clrenv and keepenv statements (see Section 4.4.4 [Environment], page 25).

4.4.1.4 User-defined variables

In addition to the built-in variables, arbitrary variables can be defined and used in the configuration file. These user-defined variables are defined using the set statement (see Section 4.4.3.1 [set], page 21) and are normally used to pass information between rules. They are invisible to whatever command rush executes as the final result of processing.

4.4.1.5 Variable Expansion

Most statements in the configuration file undergo variable expansion prior to their use. During variable expansion, references to variables in the string are replaced with their actual values. A variable reference has two basic forms:

      $v
      ${v}

where v is either the name of the variable (request, environment, or userdefined), or the index of the positional variable. The notation in curly braces serves several purposes. First, it is obligatory if v is an index of the positional variable that is negative or greater than 9. Secondly, it should be used if the variable reference is immediately followed by an alphanumeric symbol, which will otherwise be considered part of it (as in ‘${home}dir’). Finally, this form allows for specifying the action to take if the variable is undefined or expands to an empty value.

The following special forms are recognized:

${variable:-word} Use Default Values. If variable is unset or null, the expansion of word is substituted. Otherwise, the value of variable is substituted.

${variable:=word} Assign Default Values. If variable is unset or null, the expansion of word is assigned to variable. The value of variable is then substituted.

${variable:?word} Display Error if Null or Unset. If variable is null or unset, the expansion of word (or a message to that effect if word is not present) is output to the current logging channel. Otherwise, the value of variable is substituted.

${variable:+word} Use Alternate Value. If variable is null or unset, nothing is substituted, otherwise the expansion of word is substituted.

These constructs test for a variable that is unset or null. Omitting the colon results in a test only for a variable that is unset.

When expanding a variable reference, the variable name is first looked among the request variables. If it is not found, it is looked up in the userdefined variable list. If it is not there, the look up in the environment is attempted.

If the variable name is not found in any of these lists, the default rush behavior is to report the error of ‘config-error’ class (see Section 4.3.4 [Error Messages], page 12) and exit. To gracefully handle such cases, use the default value construct, defined above. For example, the following statement safely appends the string ‘/opt/man’ to the value of the MANPATH environment variable:

    setenv MANPATH = "${MANPATH:-""}${MANPATH:+:}/opt/man"

The ‘${MANPATH:-""}’ reference ensures no error is reported if the variable is undefined. The ‘${MANPATH:+:}’ reference appends a semicolon to the value, if the variable is defined. Finally the string ‘/opt/man’ is appended to the resulting value.

Another way to gracefully handle undefined variables, is to use the expand-undefined global setting. If you place the following statement at the beginning of your configuration file, any undefined variable will be silently expanded to empty string:

    global
      expand-undefined true

This statement affects variable expansion in statements that follow it in the configuration file. So you can place it in some point after which this behavior is needed, and then disable it where it is no longer desired, by using the following global statement:

    global
      expand-undefined false

4.4.2 Matching Conditions

[rule] match expr

The match statement defines conditions that decide whether the rule matches the particular request. Its argument is a simple expression or a boolean expression involving several simple expressions.

A simple expression is either a comparison or membership test.

4.4.2.1 Comparisons

A comparison expression is:

    lhs op rhs

here, lhs (left-hand side) is a string (quoted or unquoted), or a variable reference (see Section 4.1 [Lexical Structure], page 7), rhs (right-hand side) is a string or number, and op is one of the following binary operators:

‘==’  Equality (string or numeric)
‘!=’  Inequality (string or numeric)
‘<’  Less than
‘<=’  Less than or equal to
‘>’  Greater than
‘>=’  Greater than or equal to
‘~’  Regexp matching
‘!~’  Negated regexp matching

Table 4.4: Comparison Operators

Prior to evaluating simple expression, its left-hand side undergoes variable expansion and backreference interpretation. In contrast, the right-hand side is always treated verbatim.

For example the following rule will match any request with 2 or more arguments (recall, that the command name itself is counted as one of the arguments):

    rule
      match $# >= 2

The ‘==’ and ‘!=’ can operate both on strings and on numbers. When applied to strings the ‘==’ means byte-to-byte equality, e.g.

      match $0 == "/bin/ls"

will match requests with ‘/bin/ls’ as the command name.

The ‘~’ and ‘!~’ operators implement regular expression matching.

The expression ‘lhs ~ rx’ yields ‘true’ if lhs matches regular expression rx. E.g.

      match $command ~ "^scp (-v )?-t /incoming/(alpha|ftp)"

The ‘!~’ evaluates to ‘true’ if lhs does not match the regular expression in the rhs.

If the regular expression contains parenthesized groups, subsequent commands can refer to the strings that matched the groups using the backreference notation ‘%n’, where n is 1-based index ordinal number of the group in the regular expression (see [backreference], page 8). The reference ‘%0’ expands to the entire matched string. For example:

    rule chdir
      match $command "^cd (.+) && (.+)"
      chdir %1
      set command = %2
      fall-through

It splits the compound command into the working directory and the command itself. Then it remembers the name of the working directory (first parenthesized group – ‘%1’) for changing to it later (see [chdir], page 28) and resets the command line to the part of the string that follows the ‘&&’ token. Finally, it passes control to another rules (see Section 4.4.7 [Fall-through], page 29).

4.4.2.2 Membership operators

Membership operators check if their argument is a member of some set of values. There are two such operators.

    lhs in ( args )

The in operator evaluates to ‘true’ if lhs is listed in args, which is a whitespace-separated list of strings. For example:

      match $0 in ("scp" "rsync")

The group operator evaluates to ‘true’ if the requesting user is a member of at least one group listed in its right-hand side. It can have two forms:

group grp Evaluate to ‘true’ if the user is a member of the group grp. The group can be given either by its name or GID.

group ( list )

Evaluate to ‘true’ if the user is a member of one of the groups in whitespace delimited list. Members of list are group names or GIDs.

4.4.2.3 File system tests

File system tests check file types and ownership. They are similar to options to test shell command:

-b file file exists and is block special -c file file exists and is character special -d file file exists and is a directory -e file file exists -f file file exists and is a regular file -g file file exists and is set-group-ID -G file file exists and is owned by the primary group of the current user.

-h file

-L file file exists and is a symbolic link -k file file exists and has its sticky bit set -L file file exists and is a symbolic link (same as -h) -O file file exists and is owned by the current user -p file file exists and is a named pipe -r file file exists and read permission is granted -s file file exists and has a size greater than zero -S file file exists and is a socket -u file file exists and its set-user-ID bit is set -w file file exists and write permission is granted -x file file exists and execute (or search) permission is granted

4.4.2.4 Boolean expressions

Simple expressions can be combined into complex conditions using boolean operators:

‘||’  Disjunction (or)
‘&&’  Conjunction (and)
‘!’  Negation

Table 4.5: Boolean Operators

Arguments to these operators can be either simple expressions or another boolean expressions. The operators in the table above are ordered by their precedence. As in most programming languages, parentheses can be used to enforce the desired order of evaluation.

Both binary operators implement shortcut evaluation.

For example, the following rule will match if the command name contains

‘git-receive-pack’ or ‘git-upload-pack’ and either the UID is 100 or the

user is a member of the group ‘git’:

    rule
      match $0 ~ "git-(receive|upload)-pack" && \
                 ($uid == 100 || group "git")

Notice the use of parentheses to enforce proper evaluation order. The ‘&&’ operator has higher priority than ‘||’. Without parentheses the rule would match if either the command name matched the regexp and the user ID was 100, or if the user was a member of the ‘git’ group, no matter what command was issued.

4.4.3 Modifying variables

Rules can change or unset variables. Two separate groups of statements are provided to that effect. The set, unset, and map statements operate on positional, request, and user-defined variables. The setenv, unsetenv, clrenv, and keepenv statements modify the environment. These will be discussed in a separate subsection (see Section 4.4.4 [Environment], page 25).

Modifications to positional and request variables deserve a special explanation.

The only two request variables that can be modified (but not unset) are

$command and $program.

Positional variables and the $command request variable are mutually dependent. If the $command is modified, the word splitting is applied to it and resulting words are assigned to the positional variables. Similarly, any modifications to positional variables trigger rebuilding of the $command variable from the modified arguments. Both operations are run immediately after the change that triggered them. Notice, however, that any transformations, including variable modifications, are executed after match statements have been evaluated, so that match always operates on unchanged variables, no matter where in the rule you place it,

If the rules result in accepting the request, then modified $command becomes the actual command that rush will execute.

Obviously, none of the request variables can be unset. You can however, unset a positional variable (excepting ‘$0’). It is equivalent to removing the corresponding argument from the command line.

4.4.3.1 The set statement

The set statement modifies the value of a positional, request, or user-defined variable.

[rule] set name = value
[rule] set [n] = value

Sets the variable name to value. Prior to use, value undergoes backreference interpretation (see [backreference], page 8) and variable expansion (see Section 4.4.1.5 [Variable expansion], page 16). The second form assigns to the positional variable ‘$n’. It is discussed in more detail in Section 4.4.5 [Transformations], page 26.

[rule] set name = value ~ s-expr
[rule] set [n] = value ~ s-expr

Applies the sed search-and-replace expression s-expr to value and assigns the result to the variable name or argument n. Both value and s-expr are subject to variable expansion and backreference interpretation.

[rule] set name =~ s-expr
[rule] set [n] =~ s-expr

Applies the sed-like search-and-replace expression s-expr to the current value of the variable name and stores the resulting string as its new value. Prior to use, s-expr undergoes backreference interpretation (see [backreference], page 8) and variable expansion (see Section 4.4.1.5 [Variable expansion], page 16). This is a shortcut for

      set name = ${name:-""} ~ s-expr

Second form modifies the value of the positional variable ‘$n’. This statement is a shortcut for

      set [n] = ${n:-""} ~ s-expr

See Section 4.4.5 [Transformations], page 26, for a detailed discussion.

The transformation expression, s-expr, is sed-like replace expression of the form:

    s/regexp/replace/[flags]

where regexp is a regular expression, replace is a replacement for each part of the input that matches regexp and flags are optional flags that control the substitution. Both regexp and replace are described in Section “The ‘s’ Command” in GNU sed.

As in sed, you can give several replace expressions, separated by semicolons.

Supported flags are:

‘g’ Apply the replacement to all matches to the regexp, not just the first.

‘i’ Use case-insensitive matching

‘x’ regexp is an extended regular expression (see Section “Extended regular expressions” in GNU sed).

‘number’ Only replace the numberth match of the regexp.

Note: the POSIX standard does not specify what should happen when you mix the ‘g’ and number modifiers. Rush follows the GNU sed implementation in this regard, so the interaction is defined to be: ignore matches before the numberth, and then match and replace all matches from the numberth on.

Normally, the s-expr is a quoted string, and as such it is subject to backslash interpretation. It is therefore important to properly escape backslashes, especially in replace part. Consider this example:

    set bindir = $program ~ "s/(.*)\\//\\1/"

The intention is to extract the directory part of the executable program name and store it in the variable ‘bindir’. Notice, that each backslash is escaped, so that the actual string that is compiled into a regular expression is

    s/(.*)\//\1/

4.4.3.2 The insert statement

The insert statement inserts new positional argument at a given position. Its syntax is similar to set:

[rule] insert [n] = value
[rule] insert [n] = value ~ s-expr

Shift arguments starting from n one position to the right (so that n becomes n+1 etc.) and insert value at argv[n]. In the second form, the value to be inserted is computed by applying sed-expression s-expr to value. Both value and s-expr are subject to variable expansion and backreference interpretation.

Example using this statement to insert the --root=/tmp argument at position 1:

    insert [1] = "--root=/tmp"

Note that when inserting multiple arguments (e.g. an option with a value), you have two possibilities. First, you can insert each argument at its corresponding position. For example, to insert two arguments ‘--root’ and ‘/tmp’ starting at position 1, one can use:

    insert [1] = "--root"
    insert [2] = "/tmp"

Otherwise, you can revert the arguments and insert them at the same position, as shown in the example below:

    insert [1] = "/tmp"
    insert [1] = "--root"

4.4.3.3 The unset statement

[rule] unset name

Unset the variable name.

[rule] unset n

Unset the positional argument n (an integer number greater than 0), shifting the remaining arguments one position left. The effect is the same as from delete (see Section 4.4.3.5 [delete], page 24).

4.4.3.4 The remopt statement

The remopt statement removes from the command line all occurrences of the supplied option.

[rule] remopt sopt
[rule] remopt sopt lopt

Remove from the command line all occurrences of the short option described by sopt. The sopt argument is the short option letter, optionally followed by a colon if that option takes a mandatory argument, or by two colons if it takes an optional argument. Optional lopt supplies a long option equivalent to sopt. If no short option equivalent exists, use ‘_’ as sopt, eventually followed by ‘:’ or ‘::’.

For example, to remove all occurrences of the -r (--root) option that takes a mandatory argument, use:

 remopt r: root

4.4.3.5 The delete statement

Another statement modifying the command line is delete:

[rule] delete n

Delete nth argument.

[rule] delete i j

Delete positional parameters between ‘$i’ and ‘$j’, inclusive.

Neither form can be used to delete the program name (‘$0’). For example, the following statement deletes all arguments from the com-

mand line, except for the program name:

 delete 1 -1

To delete a single argument, unset can also be used. The following statements have the same effect:

 delete 2
 unset 2

4.4.3.6 The map statement

[rule] map name file delim key kn vn
[rule] map [n] file delim key kn vn default

The ‘map’ statement uses file lookup to find a new value for the variable name (or, in its second form, for the positional variable ‘$n’). Arguments are:

file Name of the map file. It must begin with ‘/’ or ‘~/’. Before using, the file permissions and ownership are checked using the procedure described in [security checks], page 7. delim A string containing allowed field delimiters.

key The value of the lookup key. Before using, it undergoes backslash interpretation and variable expansion.

kn Number of the key field in file. Fields are numbered starting from 1.

vn Number of the value field.

default If supplied, this value is used as a replacement value, when the key was not found in file.

The map file consists of records, separated by newline characters (in other words, a record occupies one line). Each record consists of fields, separated by delimiters listed in delim argument. If delim contains a space character, then fields may be delimited by any amount of whitespace characters (spaces and/or tabulations). Otherwise, exactly one delimiter delimits fields.

Fields are numbered starting from 1.

The map action works as follows:

1. Variable expansion is performed on the key argument (see Section 4.4.1.5 [Variable expansion], page 16) and the resulting value is used as lookup key.

2. The file is scanned for a record whose knth field matches the lookup key.

3. If such a record is found, the value of its vnth field is assigned to the variable.

4. Otherwise, if default is supplied, it becomes the new value of the variable.

5. Otherwise, the variable remains unchanged.

For example, suppose that the file /etc/passwd.rush has the same syntax as the system passwd file (see Section “passwd” in passwd(5) man page). Then, the following statement will replace ‘$0’ with the value of ‘shell’ field, using the current user name as a key:

    map [0] /etc/passwd.rush : ${user} 1 7

See also Section 4.4.11 [Interactive], page 33, for another example of using this statement.

4.4.4 Environment

The following actions modify the environment in which the program will be executed.

[rule] clrenv

Clear the environment.

[rule] keepenv list

Retain the names in list in the environment. This statement should be used in conjunction with clrenv. Argument is a whitespace delimited list of variables to retain. Each element in the list can be either a variable name, or a shell-style globbing pattern, in which case all variables matching that pattern will be retained, or a variable name followed by an equals sign and a value, in which case it will be retained only if its actual value equals the supplied one. For example, to retain only variables with names beginning with ‘LC_’:

    keepenv "LC_*"
[rule] setenv name = value

Set the environment variable name. The value argument is subject to variable expansion (see Section 4.4.1.5 [Variable expansion], page 16) and backreference interpretation (see [backreference], page 8). For example, to modify the PATH value:

    setenv PATH = "$PATH:/opt/bin"
[rule] unsetenv list

Unset environment variables listed as arguments. Argument is a whitespace delimited list of variables to retain. Each element in the list can be either a variable name, or a shell-style globbing pattern, in which case all variables matching that pattern will be unset, or a variable name followed by an equals sign and a value, in which case it will be unset only if its actual value equals the supplied one.

[rule] evalenv string

Performs backslash interpretation, backreference interpretation and variable expansion on string and discards the result. This statement is similar to the shell’s colon statement. For example, the following statement will define the DEPTH variable and initialize it to 10, unless it is already defined:

    evalenv ${DEPTH:=10}

4.4.5 Transformations

Transformations are special actions that modify entire command line or particular arguments from it (positional variables).

Statements that modify variable have been described in the previous section: these are set, insert, unset, remopt, delete and map statements. When set or map is applied to the ‘command’ variable, it modifies entire command line. When these statements are applied to an index (‘[n]’), they modify the corresponding positional variable (argument). This subsection discusses the implications of modifying these variable and illustrates them with some examples.

Positional variables and the $command request variable are mutually dependent. If the $command is modified, the word splitting is applied to it and resulting words are assigned to the positional variables. Similarly, any modifications to positional variables trigger rebuilding of the $command variable from the modified arguments. See Section 4.4.3 [Modifying variables], page 21, for more detail on it.

Let’s consider several examples.

1. Echo the command line

rule
set command = "/bin/echo $command"

2. Remove all occurrences of -r option and its arguments from the command line, and then adds its own -r option and replaces ‘svnserve’ with the full program file name. There are at least three different ways to do so. a. The recommended approach is to use the remopt and insert state-

ments, as shown below:

rule svn
match $command ~ "^svnserve -t"
set program = "/usr/bin/svnserve"
remopt r:
insert [1] = "-r"
insert [2] = "/svnroot"

b. The same can be achieved using regular expressions. This was the

default in versions of rush prior to 2.0:

rule svn
match $command ~ "^svnserve -t"
set command =~ "s/-r *[^]*//"
set command =~ \
    "s|^svnserve |/usr/bin/svnserve -r /svnroot |"

Notice the use of ‘|’ as a delimiter in s-command, in order to avoid escaping each ‘/’ in the pathname. Without it, the expression in the second set command will be

"s/^svnserve /\\/usr\\/bin\\/svnserve -r \\/svnroot /"

c. The same rule, rewritten using the single set statement:

rule svn
match $command ~ "^svnserve -t"
set command =~ "s|-r *[^]*||;\
       s|^svnserve |/usr/bin/svnserve -r /svnroot |"

3. Override the executable program name.

rule cvs
match $command ~ "^cvs server"
set [0] = /usr/bin/cvs

4.4.6 System Actions

System actions provide an interface to the operating system.

[rule] umask mask

Set the umask. The mask must be an octal value not greater than ‘0777’. The default umask is ‘022’.

[rule] newgrp group-id
[rule] newgroup group-id

Change the current group ID to group-id, which is either a numeric value or a name of an existing group.

[rule] chroot dir

Change the root directory to that specified in dir. This directory will be used for file names beginning with ‘/’. The argument is subject to tilde, variable, and backreference expansions. During tilde expansion, a tilde (‘~’) at the start of string is replaced with the absolute pathname of the user’s home directory. The two other expansions are described in Section 4.4.1.5 [Variable expansion], page 16, and [backreference], page 8.

The directory dir must be properly set up to execute the commands. For example, the following rule defines execution of sftp-server in an environment chrooted to the user’s home directory:

 rule sftp
   match $program ~ "^.*/sftp-server"
   set [0] = "bin/sftp-server"
   chroot "~"

For this to work, each user’s home must contain the directory bin with a copy of sftp-server in it, as well as all directories and files that are needed for executing it, in particular lib.

[rule] chdir dir

Change to the directory dir. The argument is subject to tilde, variable (see Section 4.4.1.5 [Variable expansion], page 16), and backreference expansions (see [backreference], page 8). If both chdir and chroot are specified, then chroot is applied first.

[rule] limits res

Impose limits on system resources, as defined by res. The argument consists of commands, optionally separated by any amount of whitespace. A command is a single command letter followed by a number, that specifies the limit. The command letters are case-insensitive and coincide with those used by the shell ulimit utility:

Command  The limit it sets
A  max address space (KB)
C  max core file size (KB)
D  max data size (KB)
F  maximum file size (KB)
M  max locked-in-memory address space (KB)

N  max number of open files
R  max resident set size (KB)
S  max stack size (KB)
T  max CPU time (MIN)
U  max number of processes

L max number of logins for this user (see below)

P process priority -20..20 (negative = high priority)

For example:

    limits T10 R20 U16 P20

If some limit cannot be set, execution of the rule aborts. In particular, the ‘L’ limit can be regarded as a condition, rather than an action. Setting limit Ln succeeds only if no more than n rush instances are simultaneously running for the same user. This can be used to limit the number of simultaneously open sessions. The use of ‘L’ resource automatically enables forked mode. See Section 4.4.8 [Accounting and Forked Mode], page 30, for more information about it.

4.4.7 Fall-through

The fall-through statement is a special action that does not execute the requested command. When a matching fall-through rule is encountered, rush evaluates it and continues scanning its configuration for the next matching rule. Any modifications to the request found in the fall-through rule take effect immediately, which means that subsequent rules will see modified command line and environment. Execution of any other actions found in the fall-through rule is delayed until a usual rule is found.

A fall-through rule is declared using the following statement:

[rule] fall-through
[rule] fallthrough

Declare a fall-through rule.

Usually this statement is placed as the last statement in a rule, e.g.:

 rule default
   umask 002
   clrenv
   keepenv HOME USERNAME PATH
   fall-through

Fall-through rules provide a way to set default values for subsequent rules. For example, any rules that follow the ‘default’ rule shown above, will inherit the umask and environment set there.

One can also use fall-through rules to “normalize” command lines. For example, consider this rule:

    rule default
      set [0] =~ "s|.*/||"
      fall-through

It will remove all path components from the first command line argument. As a result, all subsequent rules may expect a bare binary name as the first argument.

Yet another common use for such rules is to enable accounting (see the next subsection), or set resource limits for the rest of rules:

    rule default
      limit l1
      fall-through

4.4.8 Accounting and Forked Mode

GNU Rush is able to operate in two modes, which we call default and forked. When operating in the default mode, the process image of rush itself is overwritten by the command being executed. Thus, when it comes to launching the requested command, the running instance of rush ceases to exist.

There is also another operation mode, which we call forked mode. When running in this mode, rush executes the requested command in a subprocess, and remains in memory supervising its execution. Once the command terminates, rush exits.

One advantage of the forked mode is that it allows you to keep accounting, i.e. to note who is doing what and to keep a history of invocations. The accounting, in turn, can be used to limit simultaneous executions of commands (logins, in GNU Rush terminology), as requested by ‘L’ command to limit statement (see [L limit], page 29).

The forked mode is enabled on a per-rule basis, for rules that contain either ‘L’ command in the limit statement, or ‘acct on’ command:

[rule] acct bool

Turn accounting mode on or off, depending on bool. The argument can be one of the following: ‘yes’, ‘on’, ‘t’, ‘true’, or ‘1’, to enable accounting, and ‘no’, ‘off’, ‘nil’, ‘false’, ‘0’, to disable it. Notice, that there is no need in explicit acct on command, if you use

  limit L.

The notion ‘rule contains’, used above, means that either the rule in question contains that statement, or inherits it from one of the fall-through rules (see Section 4.4.7 [Fall-through], page 29) that were matched before it. In fact, in most cases the accounting should affect all rules, therefore we suggest to enable it in a fall-through rule at the beginning of the configuration file, e.g.:

    rule default
      acct on
      fall-through

If the need be, you can disable it for some of the subsequent rules by placing acct off in it. Notice, that this will disable accounting only, the forked mode will remain in action. To disable it as well and enforce default mode for a given rule, use the following statement:

[rule] fork bool

Enable or disable forked mode. This statement is mainly designed as a way of disabling the forked mode for a given rule.

Once accounting is enabled, you can use the rushwho command to see the list of users presently running some commands (see Chapter 9 [Rushwho], page 51) and view the history of last accesses using rushlast command (see Chapter 10 [Rushlast], page 55).

4.4.9 Post-process Notification

Rush can be configured to send a notification over INET or UNIX sockets, after completing user request. It is done using the post-socket statement:

[rule] post-socket url

Notify URL about completing the user request. This statement implies forked mode (see Section 4.4.8 [Accounting and Forked Mode], page 30). Allowed formats for url are: inet://hostname[:port]

Connect to remote host hostname using TCP/IP. Hostname is the host name or IP address of the remote machine. Optional port specifies the port number to connect to. It can be either a decimal port number or a service name from /etc/services. If port is absent, ‘tcpmux’ (port 1) is assumed.

unix://filename

local://filename Connect to a UNIX socket filename.

For example:

      rule default
        post-socket "inet://localhost"

The GNU Rush notification protocol is based on TCPMUX (RFC 1078

(http://www.rfc-editor.org/rfc/rfc1078.txt)).

After establishing connection, rush sends the rule tag followed by a CRLF pair. The rule tag acts as a service name. The remote party replies with a single character indicating positive (‘+’) or negative (‘-’) acknowledgment, optionally followed by a message of explanation, and terminated with a CRLF.

If positive acknowledgment is received, rush sends a single line, consisting of the user name and the executed command line, separated by a single space character. The line is terminated with a CRLF.

After sending this line, rush closes the connection.

The post-process notification feature can be used to schedule execution of some actions after certain rules.

See Section 6.7 [notification example], page 45, for an example of how to use this feature.

4.4.10 Exit rule

The exit rule does not execute any commands. Instead, it writes the supplied error message to the specified file descriptor and exits immediately. The exit rule is defined using the following statement:

[rule] exit fd message
[rule] exit message

Write textual message message to a file descriptor, given by the optional argument fd. If fd is absent, ‘2’ (standard error) is used.

The message argument can be either a quoted string, or an identifier. If it is a quoted string, it is subject to backreference interpretation and

variable expansion prior to being used.

For example (note the use of line continuation character):

    exit "\
        \r\nYou are not allowed to execute that command.\r\n\
        \r\nIf you think this is wrong, ask <foo@bar.com> for assistance.\r\n"

If message is an identifier, it must be the name of a predefined error message (see Section 4.3.4 [Error Messages], page 12). The corresponding message text will be printed. For example:

      exit nologin-message

If the identifier does not match any predefined error message name, an error of type ‘config-error’ is signaled and rush exits.

Exit actions are useful for writing trap rules, i.e. the rules that are intended to trap incorrect or prohibited command lines and to return customized reply messages in such cases. Consider the following rule:

    rule git
      match $program ~ "^git-.+" && $1 ~ "^/sources/[^]+\.git$"
      set command =~ "s|.*|/usr/bin/git-shell -c \"&\"|"

It allows the client to use only those Git repositories that are located under /sources directory^{3}. If a user tries to access a repository outside this root, he will be returned a default error message, saying ‘You are not permitted to execute this command’ (see Section 4.3.4 [Error Messages], page 12). You can, however, provide a more convenient message in this case. To do so, place the following after the ‘git’ rule:

3 See Section 6.6 [git], page 44, for a better way to handle Git accesses.

    rule git-trap
      match $command ~ "^git-.+"
      exit "fatal: Use of this repository is prohibited."

This rule will trap all git invocations that do not match the ‘git’ rule.

4.4.11 Interactive Access

Sometimes it may be necessary to allow some group of users limited access to interactive shells. GNU Rush contains provisions for such usage. When rush is invoked without -c it assumes interactive usage. In this case only rules explicitly marked as interactive are considered, the rest of rules is ignored.

[rule] interactive bool

If bool is ‘true’, this statement marks the rule it appears in as interactive. This rule will match only if rush is invoked without command line arguments.

Unless command line transformations are applied, interactive rule finishes by executing /bin/sh. The first word in the command line (argv[0]) is normally set to the base name of the command being executed prefixed by a dash sign. Consider the following example:

    rule login
      interactive true
      group rshell
      map program /etc/rush.shell : ${user} 1 2
      set [0] = ${program} ~ "s|^.*/||;s,^,-r,"
    rule nologin
      interactive true
      exit You don't have interactive access to this machine.

The ‘login’ rule will match interactive user requests if the user is a member of the group ‘rshell’. It uses /etc/rush.shell to select a shell to use for that user (see Section 4.4.3.6 [map], page 24). This map file consists of two fields, separated by a colon. If the shell is found, its base name, prefixed with ‘-r’, will be used as ‘argv[0]’ (this indicates a restricted login shell). Otherwise, the trap rule ‘nologin’ will be matched, which will output the given diagnostics message and terminate rush.

To test interactive access, use the -i option:

    rush --test -i

4.4.12 Localization

GNU Rush is internationalized, which means that it is able to produce log and diagnostic messages in any language, if a corresponding translation file is provided. This file is called a localization or domain file. To find an appropriate localization file, rush uses the following parameters: locale Locale name is a string that describes the language, territory

and optionally, the character set to use. It consists of the language (ISO 639) and country (ISO 3166) codes, separated by an underscore character, e.g. ‘en_US’ or ‘pl_PL’. If a character set is specified, its name follows the country code and is separated from it by a ‘@’ character. There are two special locale names: ‘C’ or ‘POSIX’ mean to use the default POSIX locale, and ‘""’ (an empty string), means to use the value of the environment variable LC_ALL as the locale name.

locale dir Directory where localization files are located. If not specified, a predefined set of directories is searched for the matching file.

domain Text domain defines the base name of the localization file.

Given these parameters, the name of the full pathname of the localization file is defined as:

    locale_dir/locale/LC_MESSAGES/domain.mo

GNU Rush produces three kinds of messages:

diagnostics These are diagnostics messages that GNU Rush produces to its log output (syslog, unless in test mode).

error messages Messages sent to the remote party when rush is not able to execute the request (see Section 4.3.4 [Error Messages], page 12).

exit messages These are messages sent to the remote party by exit rules (see Section 4.4.10 [Exit], page 32).

These messages use different domain names (and may use different locale directories). The diagnostics and error messages use textual domain ‘rush’. The corresponding locale directory is defined at compile time and defaults to prefix/share/locale, where prefix stands for the installation prefix, which is /usr/local, by default.

GNU Rush is shipped with several localization files, which are installed by default. As of version 2.2, these files cover the following languages: Chinese, Danish, Dutch, Finnish, French, Galician, German, Polish, Portuguese, Serbian, Spanish, Swedish, Ukrainian, and Vietnamese. If the localization you need is not in this list, visit

http://translationproject.org/domain/rush.html. If it is not there

either, consider writing it (see Section “Translators” in GNU gettext utilities, for a detailed instructions on how to do that).

Exit messages use custom domain files. It is the responsibility of the system administrator to provide and install such files.

4.4.12.1 Localization Directives

The following configuration directives control localization. They are available for use in rule statements:

[rule] locale name

Sets the locale name. To specify empty locale, use ‘""’ as name (recall that empty locale name means to use the value of the environment variable LC_ALL as locale name).

[rule] locale-dir name

Sets the name of the locale directory.

[rule] text-domain name

Sets the textual domain name.

The following configuration fragment illustrates their use:

rule l10n
  locale "pl_PL"
  text-domain "rush-config"
  fall-through

Different users may have different localization preferences. See [per-user l10n], page 36, for a description of how to implement this.

4.4.12.2 Writing Your Localization

You need to write a localization file for your configuration script if it implements exit rules (see Section 4.4.10 [Exit], page 32) and changes user locale (see Section 4.4.12.1 [Localization Directives], page 35).

Preparing a localization consists of three stages: extracting exit messages and forming a PO file, editing this file, compiling and installing it. The discussion below describes these stages in detail.

1. Creating a ‘po’ file. A PO (Portable Object) file is a plain text file, containing original messages and their translations for a particular language. See Section “PO Files” in GNU gettext utilities, for a description of its format. The script rush-po extracts translatable messages from the configuration file and produces a valid PO file. It takes the name of the rush configuration file as its argument and produces the PO file on the standard output, or in the file given with the -o (--output) option. E.g., to create a PO file from your configuration file, run:

    rush-po -o myconf.po /usr/local/etc/rush.rc

2. Editing the PO file Open the created PO file with your favorite editor and supply message translations after msgstr keywords. Although you can use any editor capable of handling plain text files, we recommend to use GNU Emacs, which provides a special po-mode. See Section “Basics” in GNU gettext utilities, for guidelines on editing PO files and using the po-mode.

3. Compiling the PO file When ready, the PO file needs be compiled into a MO (Message Object) file, which is directly readable by rush. This is done using msgfmt utility from GNU gettext:

          msgfmt -o myconf.mo myconf.po

See Section “msgfmt Invocation” in GNU gettext utilities, for a detailed description of the msgfmt utility. After creating the MO file, copy it into appropriate directory. It is important that the installed MO file uses the naming scheme described in [localization file naming], page 34.

4.5 Include

The include statement forces inclusion of the named file in that file location:

[rule] include file

Include file file.

The statement is evaluated when parsing the configuration file, which means that file undergoes only tilde expansion: the two characters ‘~/’ appearing at the beginning of file are replaced with the full path name of the current user’s home directory.

If file is a directory, that directory is searched for a file whose name coincides with the current user name. If such a file is found, it is included.

In any case, if the file named by file (after tilde expansion) does not exist, no error is reported, and parsing of the configuration file continues.

Before including the file rush checks if it is secure, using the same rules as for the main configuration file (see [security checks], page 7). The exact list of checks can be tuned using the include-security statement (see Section 4.3.6 [include-security], page 13).

The include statement can be used only within a rule. The included file may not contain rule and global statements.

This statement provides a convenient way for user-dependent rush configuration. For example, the following fall-through rule (see Section 4.4.7 [Fall-through], page 29) allows the administrator to keep each user personal configuration in a file named .rush, located in the user’s home directory:

    rule user
      include "~/.rush"
      fall-through

Of course, it is supposed that such a per-user file, if it exists, is writable only for super-user.

The use of include files may be especially useful for per-user localization (see Section 4.4.12 [Localization], page 33). It suffices to provide a fallthrough rule, similar to the one above, and to place a locale directive in ~/.rush files, according to the user preferences.

5 Default Configuration

You can compile rush with the default configuration built in the binary. Such a binary can then be run without configuration file. However, if a configuration file is present, it will be used instead of the built-in configuration.

To compile rush with the built-in configuration, first compile the package as usual. Then, prepare a configuration file, and test it using rush --lint. If the test shows no errors, reconfigure the package, using the

--with-default-config option:
      ./configure --with-default-config=file

where file is the name of your configuration file. Then, recompile and install the package.

You can inspect the built-in configuration using the --show-default option:

      rush --show-default

6 Usage Tips

In this chapter we will explain how to write GNU Rush configuration rules for several popular remote copy and version control system utilities. For this purpose, we assume the following setup:

• Users are allowed to use scp and rsync to upload files to the /incoming directory and to copy files to and from their ~/public_html directory. The actual location of the /incoming directory is /home/ftp, but that must be transparent to users, i.e. they use scp file host:/incoming

(not host:/home/ftp/incoming) to upload files.

• Additionally, users may use sftp to manage their ~/public_html directory. In this case, to prevent users from accessing other directories, sftp-server is executed in a chrooted environment.

• The server runs three version control system repositories, whose corresponding root directories are:

VCS  Repository Root
cvs  /cvsroot
svn  /svnroot
git  /gitroot

6.1 scp

The scp utility is executed on the server side with option -t, when copying files to server, and with -f when copying from it. Thus, the basic templates for scp rules are:

# Copying to server:
rule scp-to
  match $command ~ "^scp -t"
  ...
# Copying from server:
rule scp-from
  match $command ~ "^scp -f"
  ...

You may also wish to allow for -v (‘verbose’) command line option. In this case, the ‘scp-to’ rule will become:

rule scp-to
  match $command ~ "^scp (-v )?-t"
  ...

Now, we want users to be able to upload files to /home/ftp/incoming directory. Moreover, the /home/ftp directory prefix must be invisible to them. We should also make sure that the user cannot get outside the incoming directory by using ../ components in his upload path. So, our first rule for scp uploads will be:

    rule scp-to-incoming
      match $command ~ "^scp (-v )?-t /incoming/" && \
                       ${-1} !~ "\\.\\./"
      set command "/bin/scp"
      set [-1] =~ "s|^|/home/ftp/|"

The match statement ensures that no relative components are used. The two set statements ensure that the right scp binary is used and that /home/ftp prefix is prepended to the upload path.

Other than uploading to /incoming, users must be able to use scp to manage public_html directories located in their homes. They should use relative paths for that, i.e., the command:

    $ scp file.html server:
will copy file file.html to ~/public_html/file.html on the server. The

corresponding rule is:

    rule scp-home
      match $command ~ "^scp (-v )?-[tf] [^/].*" && \
                       ${-1} !~ "\\.\\./"
      set [0] = "/bin/scp"
      set [-1] =~ "s|^|public_html/|"
      chdir "~"

Finally, we provide two trap rules for diagnostic purposes:

    rule scp-to-trap
      match $command ~ "^scp (-v )?-t"
      exit "Error: Uploads to this directory prohibited"
    rule scp-from
      match $command ~ "^scp (-v )?-f"
      exit Error: Downloads from this directory prohibited

6.2 rsync

On the server side, rsync is executed with the --server command line option. In addition, when copying files from the server, the --sender option is used. This makes it possible to discern between incoming and outgoing requests.

In our setup, rsync is used the same way as scp, so the two rules will be:

rule rsync-incoming
match $command ~ "^rsync --server" && \
      $command !~ --sender && \
      ${-1} ~ "/incoming/" && ${-1} !~ "\\.\\./"
set [0] =~ "s|^|/usr/bin/|"
set [-1] =~ "s|^|/home/ftp/|"
rule rsync-home
match $command ~ "^rsync" && \
      ${-1} !~ "^[^/]" && \
      ${-1} !~ "\\.\\./"
set [0] = "s|^|/usr/bin/|"
set [-1] =~ "s|^|public_html/|"
chdir "~"

The trap rules for rsync are trivial:

rule rsync-to-trap
match $command ~ "^rsync.*--sender"
exit "Error: Downloads from this directory prohibited"
rule rsync-from-trap
match $command ~ "^rsync"
exit "Error: Uploads to this directory prohibited"

6.3 sftp

Executing sftp on the client machine invokes sftp-server, without arguments, on the server.

We want to allow our users to use sftp to manage their public_html directories. The sftp-server will be executed with the user’s home directory as root, in a chrooted environment. For this to work, each user’s home must contain a copy of sftp-server (which we’ll place in ~/bin subdirectory) and all files it needs for normal execution: /etc/group and /etc/passwd with one entry (for the user and his group), and, unless the binary is linked statically, all the shared libraries it is linked with, in the subdirectory ~/lib.

Given these prerequisites, the following rule will ensure proper sftp interaction:

rule sftp-incoming
match $command ~ "^.*/sftp-server"
set [0] = "/bin/sftp-server"
chroot "~"
chdir "public_html"

Notice the last action. Due to it, users don’t have to type cd public_html at the beginning of their sftp sessions.

6.4 cvs

Using cvs over ssh invokes cvs server on the server machine. In the simplest case, the following rule will do to give users access to CVS repositories:

    rule cvs
      match $command ~ "^cvs server"
      set command ~ "s|^cvs|/usr/bin/cvs -f"

However, cvs as of version 1.12.13 does not allow to limit root directories that users are allowed to access. It does have --allow-root option, but unfortunately this option is ignored when invoked as cvs server. To restrict possible roots, we have to run cvs in a chrooted environment. Let’s suppose we created an environment for cvs in directory /var/cvs, with the cvs binary located in /var/cvs/bin and repository root directory being /var/cvs/cvsroot. Then, we can use the following rule:

    rule cvs
      match $command ~ "^cvs server"
      set [0] = "/bin/cvs"
      chroot "/var/cvs"

6.5 svn

Remote access to SVN repositories is done via svnserve binary. It is executed on server with -t option. The -r option can be used to restrict access to a subset of root directories. So, we can use the following rule:

    rule svn
      match $command ~ "^svnserve -t"
      set command =~ "s|-r *[^]*||"
      set command =~ \
          "s|^svnserve |/usr/bin/svnserve -r /svnroot|"

The first set command action removes any -r options the user might have specified and enforces a single root directory. A more restrictive action can be used to improve security:

      set command =~ "s|.*|/usr/bin/svnserve -r /svnroot|"

6.6 git

Remote access to Git repositories over ssh causes execution of git-receive-pack and git-upload-pack on the server. The simplest rule for Git is:

    rule git
      set $command ~ "^git-(receive|upload)-pack"
      set [0] =~ "s|^|/usr/bin/|"

The set action is necessary to ensure the proper location of Git binaries to use. This example supposes they are placed in /usr/bin, you will have to tailor it if they are located elsewhere on your system.

To limit Git accesses to repositories under /gitroot directory, modify the ‘$1’, as shown in the example below:

    rule git
      match $command ~ "^git-(receive|upload)-pack"
      set [1] =~ "^/gitroot[^]+\.git$"
      set [0] =~ "s|^|/usr/bin/|"

To provide more helpful error messages, you may follow this rule by a trap rule (see Section 4.4.10 [Exit], page 32):

# Trap the rest of Git requests:

    rule git-trap
      match $command ~ "^git-.+"
      exit "fatal: access to this repository is denied."

6.7 Notification

In this section we will show how to set up a mail notification for Rush rules. Let’s suppose we wish to receive emails for each upload by scp-to rule (see Section 6.1 [scp], page 41). To do so, we add the following fall through rule to the beginning of rush.rc:

    rule default
      post-socket "inet://localhost"
      fall-trough

This will enable notifications for each rule located below this one. Missing port in post-socket statement means rush will be using the default ‘tcpmux’ port.

To receive and process these requests, you will need an inetd capable to handle TCPMUX. We recommend the one from GNU Inetutils

package (GNU Inetutils (http://www.gnu.org/software/inetutils)). In
/etc/inetd.conf file, we add:

# Enable TCPMUX handling.

    tcpmux          stream  tcp  nowait root  internal

# Handle ‘scp-to’ service.

    tcpmux/+scp-to  stream  tcp  nowait root  \
                            /usr/sbin/tcpd  /bin/rushmail

The program /bin/rushmail does the actual notification. Following is its simplest implementation:

#! /bin/sh
read user command
/usr/sbin/sendmail -oi -t <<EOT
From: GNU Rush Notification <devnull@localhost>
To: <root@localhost>
Subject: GNU Rush notification
Be informed that $user executed $command.
EOT

7 Test Mode

GNU Rush provides a special test mode, intended to test configuration files and to emulate execution of commands. Test mode is enabled by the --test command line option (aliases: --lint, -t). When rush is given this option, the following occurs:

1. All diagnostic messages are redirected to standard error, instead of syslog.

2. If a single non-option argument is present, it is taken as a name of the configuration file to use.

3. The configuration file is parsed. If parsing fails, the program exits with the code 1.

4. If the -c option is present, rush processes its argument as usual (see Chapter 2 [Operation], page 3), except that the command itself is not executed.

5. Otherwise, if -i option is present, rush emulates interactive usage, but does not execute the final command. An exit status of 0 means no errors, 1 means an error has occurred. You can also emulate access by a particular user, by supplying his user

name via the --user (-u) option. This option implies --test.

In test mode, you can set debugging level (see Section 4.3.2 [Debugging], page 11) from the command line, using the --debug (-d) command line option. It expects a single number specifying debugging level as its argument. The debugging level set this way overrides settings from the configuration file.

Here are several examples that illustrate the use of test mode in various cases:

1. Test default configuration file:

      $ rush --test

2. Test configuration file sample.rc:

      $ rush --test sample.rc

3. Test interactive access

      $ rush --test -i sample.rc

4. Test the configuration file and emulate execution of the command cvs

server. Use debugging level 2:

      $ rush --test --debug=2 -c "cvs server"

5. Same, but for user ‘jeff’:

      $ rush --user=jeff --debug=2 -c "cvs server"

Note, that you don’t need to specify --test along with --user or -i options.

6. Same, but use sample.rc instead of the default configuration file:

      $ rush --test --debug=2 -c "cvs server" sample.rc

7.1 Dump Mode

Dump mode is similar to test mode. The main difference is that in this mode, rush dumps to the standard output a description of the user request after performing all checks and transformations.

The mode is requested by the --dump=attr (-D attr) option. The argument attr is a comma-separated list of the names of attributes to be included in the dump, or the word ‘all’, standing for all attributes.

Additional options and arguments are the same as for the --test option. The description is formatted as a JSON object^{1} with the following at-

tributes. These are also the allowed values for the attr list: cmdline Command line after transformations. argv Array of command line arguments after transformations. prog Name of the program to be executed. If ‘null’, argv[0] will be

used.

interactive ‘0’ for normal requests, ‘1’ for interactive requests. pw name Name of the user from the system user database. pw uid UID of the user. pw gid GID of the user. pw dir Home directory of the user, as set in the system user database. umask Value of the umask (octal). chroot dir Chroot directory. home dir Current working directory. gid New GID as set by the newgrp action, or ‘-1’ if unchanged. fork Fork mode. It is a three-state attribute: ‘0’ meaning disabled,

‘1’ meaning enabled, and ‘-1’ meaning default state.

acct Accounting mode. See ‘fork’, for a description of possible val-

ues.

text domain

Textual domain for i18n.

localedir Locale directory for i18n. locale Locale name environ Dump of the environment (array of assignments). vars Defined variables, as a JSON object.

The attribute ‘all’ stands for all attribute in the same order as listed in the table above.

1 Well, almost. It diverges from the JSON standard in that slash characters are not

escaped in string objects.

8 Option Summary

This chapter provides a short summary of rush command line options.

-c command

Specify the command to run.

-C test
--security-check=test

Configure security checks for the main configuration file. See Section 4.3.6 [include-security], page 13, for the description of test argument. See [security checks], page 7, for the discussion of the available security tests.

-d number
--debug=number

Set debugging level.

--dump=attrs

-D attrs Run in request dump mode. Argument is a comma-separated list of attribute names. See Section 7.1 [dump mode], page 48, for a detailed description of the request dump mode.

-i Emulate interactive access. See Chapter 7 [Test Mode], page 47.

--show-default

Display the default built-in configuration. See Chapter 5 [Default Configuration], page 39, for more information.

-t
--test

--lint Run in test mode. An optional argument may be used with this

option to specify alternative configuration file name, e.g.:

              $ rush --lint ./test.rc

If the -c option is also specified, rush emulates the normal processing for the command, but does not execute it.

-x

--trace Print parser traces. When used twice, print lexical scanner traces as well. This option is intended for debugging.

-T Test scanner mode. This option is used by the rush testsuite.

-u name
--user=name

Emulate access by user name. This option implies --test and is valid only when used by root and in conjunction with the -c option.

-v
--version

Display program version.

50 GNU Rush – a restricted user shell

-h

--help Display a short help message. --usage Display a concise usage summary.

9 The rushwho utility.

The rushwho utility displays a list of users who are currently using rush. The utility operates on default Rush database, which is maintained if rush runs in accounting mode (see Section 4.4.8 [Accounting and Forked Mode], page 30). The following is a sample output from rushwho:

    Login   Rule   Start     Time      PID     Command
    jeff    sftp   Sun 12:17 00:58:26  10673   bin/sftp-server

The information displayed is:

Login The login name of the user.

Rule The tag of the rule he is served under (see Section 4.4 [Rule], page 14).

Start Time when the rule began execution.

Time Duration of the session.

PID PID of the running command.

Command Command line being executed.

This format is a built-in default. It may be changed either by setting the RUSHWHO_FORMAT environment variable to the desired format string, or by using --format command line option.

9.1 Rushwho Options

This section summarizes the command line options understood by rushwho utility.

-F string
--format=string

Use string instead of the default format, described in Chapter 9 [Rushwho], page 51. See Section 9.2 [Formats], page 52, for a detailed description of the output format syntax. If string begins with a ‘@’, then this character is removed from it, and the resulting string is treated as the name of the file to read. The contents of this file is the format string. The file is read literally, except that lines beginning with ‘;’ are ignored (they can be used to introduce comments). For example, rushwho --format=@formfile reads in the contents of the file named

          formfile.
-f dir
--file=dir

Use database directory dir, instead of the default. By default, database files are located in /usr/local/var/rush.

-H
--no-header

Do not display header line.

-v
--version

Display program version.

-h

--help Display a short help message.

--usage Display a concise usage summary.

9.2 Output Formats

A format string controls the output of every record from GNU Rush accounting database. It may contain following four types of objects:

Ordinary characters These are copied to the output verbatim.

Escapes An escape is a backslash (‘\\’), followed by a single character. It is interpreted as follows:

Escape  Output
\a  Audible bell character (ASCII 7)
\b  Backspace character (ASCII 8)
\e  Escape character (ASCII 27)
\f  Form-feed character (ASCII 12)
\n  Newline character (ASCII 10)

\r

Carriage return character (ASCII 13)

\t

Horizontal tabulation character

(ASCII 9)

\v

Vertical tabulation character

(ASCII 11)

\\  A single backslash (‘\’)
\"  A double-quote.

Any escape not listed in the table above results in its second character being output.

Quoted strings Strings are delimited by single or double quotes. Within a string escape sequences are interpreted as described above.

Format specifications A format specification is a kind of function, which outputs a particular piece of information from the database record.

Each format specification starts with an opening brace and ends with a closing brace. The first word after the brace is the name of the format specification. Remaining words are positional arguments followed by keyword arguments. Both are optional. When specified, keyword arguments must follow positional ones. A keyword argument begins with a colon. For example:

(time) A single format specification.

(time 10) The same format specification with the output width limited to 10 characters.

(time 10 Duration)

The ‘time’ format specification, with the output width limited to 10 characters and ‘Duration’ as a header title.

(time 10 "Session Duration" :right :format %H:%M)

The same with two keyword arguments: ‘:right’ and ‘:format’. The latter takes the string ‘%H:%M’ as its argument. Notice the use of quoted string to preserve the whitespace.

A full list of format specifications follows.

[Format Spec] newline [count]

Causes the newline character to be output. If the optional count is supplied, that many newlines will be printed

[Format Spec] tab [num]

Advance to the next tab stop in the output stream. If optional num is present, then skip num tab stops. Each tab stop is eight characters long.

The following specifications output particular fields from the database record. They all take two positional arguments: width and title.

The first argument, width sets the maximum output length for this specification. If the number of characters actually output is less than the width, they will be padded with whitespace either to the left or to the right, depending on the presence of the :right keyword argument. If the number of characters is greater than width, they will be truncated to fit. If width is not given, the field is output as is.

The second argument, title, gives the title of this column for the heading line. By default no title is output.

Every field specification accepts at least two keyword arguments. The keyword :right may be used to request alignment to the right. This keyword is ignored if width is not given.

The keyword :empty followed by a string instructs rushwho to output that string if the resulting value for this specification would otherwise be empty.

[Format Spec] user width title [:empty repl][:right]

Print the user login name.

[Format Spec] time width title [:empty repl] [:right] [:format

      date-format]

[Format Spec] start-time width title [:empty repl] [:right]

      [:format date-format]

Date and time when the session started. The :format keyword introduces the strftime format string to be used when converting the date for printing. The default value is ‘%a %H:%M’. See Appendix A [Time and Date Formats], page 61, for a detailed description of strftime format strings.

[Format Spec] stop-time width title [:empty repl] [:right]

      [:format date-format]

Time when the command finished. This specifier is meaningful only for rushlast (see Chapter 10 [Rushlast], page 55). If the command is still running, the word ‘running’ is output.

[Format Spec] duration width title [:empty repl] [:right]

Total time of the session duration.

[Format Spec] rule width title [:right]

The tag of the rule used to serve the user. See Section 4.4 [Rule], page 14, for a detailed description of rules and tags.

[Format Spec] command width title [:empty repl] [:right]

Command line being executed.

[Format Spec] pid width title [:right]

PID of the process.

For example, the following is the default format for the rushwho utility. It is written in a form suitable for use in a file supplied with the --format=@file command line option (see [format option], page 51):

 (user 10 Login)" "
 (rule 8 Rule)" "
 (start-time 0 Start)" "
 (duration 9 Time)" "
 (pid 10 PID)" "
 (command 28 Command)

10 The rushlast utility.

The rushlast utility searches back through the GNU Rush database and displays a list of all user sessions since the database was created. By default, it displays the following information:

    Login Rule  Start     Stop      Time  Command
    gray  rsync Sun 20:43 Sun 20:43 05:57 /usr/bin/rsync /upload
    jeff  sftp  Sun 20:09 running   07:17 /bin/sftp-server

Login The login name of the user.

Rule The tag of the rule he is served under (see Section 4.4 [Rule], page 14).

Start Time when the rule began execution.

Start Time when the command finished, or the word ‘running’ if it is still running.

Time Duration of the session.

Command Command line being executed.

This format is a built-in default. It may be changed either by setting the RUSHLAST_FORMAT environment variable to the desired format string, or by using --format command line option (see Section 10.1 [Rushlast Options], page 55).

10.1 Rushlast Options

This section summarizes the command line options understood by rushlast utility.

-F string
--format=string

Use string instead of the default format, described in Chapter 9 [Rushwho], page 51. See Section 9.2 [Formats], page 52, for a detailed description of the output format syntax. To read format from a file, use --format=@filename. The file is read literally, except that lines beginning with ‘;’ are ignored (they can be used to introduce comments).

-f dir
--file=dir

Use database directory dir, instead of the default. By default, database files are located in /usr/local/var/rush.

--forward

Display entries in chronological order, instead of the reverse chronological one, which is the default.

-n number
--count=number

-number Show at most number records. The form -number is provided for compatibility with the last(1) utility.

-H
--no-header

Do not display header line.

-v
--version

Display program version.

-h

--help Display a short help message.

--usage Display a concise usage summary.

11 Accounting Database

Rush accounting database is stored in the directory localstatedir/rush, where localstatedir stands for the name of the local state directory, defined at compile time. By default, it is prefix/var, where prefix is the installation prefix, which defaults to /usr/local. Thus, the default database directory is /usr/local/var/rush. You can change this default using the --localstatedir option to configure before compiling the package. The --prefix option affects it as well.

As of version 2.2, the database consists of two files, called utmp and wtmp. The wtmp file keeps information about all user sessions, both finished and still active. The utmp file contains indices to those records in wtmp, which represent active sessions.

The wtmp grows continuously, while utmp normally grows the first day or two after enabling accounting mode, and from then on its size remains without changes. If you set up log file rotation, e.g. by using logrotate (see Section “logrotate” in logrotate man page), or a similar tool, it is safe to rotate wtmp without notifying rush. The only requirement is to truncate utmp to zero size after rotating wtmp, as shown in the following logrotate.conf snippet:

    /var/run/rush/wtmp {
        monthly
        create 0640 root svusers
        postrotate
          cat /dev/null > /var/run/rush/utmp
        endscript
    }

Accounting files are owned by ‘root’ and normally are accessible only to the owner (file mode ‘600’). You may change the default permissions using the following global configuration file statements:

[global] acct-umask mask

Set umask used when accessing accounting database files. Default value is ‘022’.

[global] acct-dir-mode mode

Set mode bits for the accounting directory. The mode argument is the mode in octal.

[global] acct-file-mode mode

Set mode bits for wtmp and utmp files.

Notice, that these statements affect file and directory modes only when the corresponding file or directory is created. Rush will not change modes of the existing files.

The following sections contain a detailed description of the structure of these two files. You may skip them, if you are not interested in technical details.

11.1 The wtmp file

The wtmp file consists of variable-size entries. It is designed so that it can easily be read in both directions.

Each record begins with a fixed-size header, which is followed by three zero-terminated strings, and the record size in size_t representation. The three strings are, in that order: the user login name, the rule tag, and the full command line.

The header has the following structure:

    struct rush_wtmp {
            size_t reclen;
            pid_t pid;
            struct timeval start;
            struct timeval stop;
            char *unused[3];
    };

where:

reclen is the length of the entire record, including the size of this header. This field is duplicated at the end of the record.

pid is the PID of the command executed for the user.

start represents the time of the beginning of the user session.

stop represents the time when the user session finished. If the session is still running, this field is filled with zeros.

unused The three pointers at the end of the structure are used internally by rush. On disk, these fields are always filled with zeros.

11.2 The utmp file

The utmp file consists of a fixed-size records of the following structure:

    struct rush_utmp {
            int status;
            off_t offset;
    };

The fields have the following meaning:

status Status of the record: ‘0’ if the record is unused, and ‘1’ if it represents an active session.

offset Offset of the corresponding record in wtmp (see previous section).

12 How to Report a Bug

Email bug reports to bug-rush@gnu.org. Please include a detailed description of the bug and information about the conditions under which it occurs, so we can reproduce it. To facilitate the task, the following list shows the basic set of information that is needed in order to find the bug:

• Package version you use.

• A detailed description of the bug.

• Conditions under which the bug appears.

• It is often helpful to send the contents of config.log file along with your bug report. This file is created after running ./configure in the GNU Rush source root directory.

Appendix A Time and Date Formats

This appendix documents the time format specifications understood by the :format keyword in time output format specifiers (see Section 9.2 [Formats], page 52). Essentially, it is a reproduction of the man page for GNU strftime function.

Ordinary characters placed in the format string are reproduced without conversion. Conversion specifiers are introduced by a ‘%’ character, and are replaced as follows: %a The abbreviated weekday name according to the current locale.

%A The full weekday name according to the current locale.

%b The abbreviated month name according to the current locale.

%B The full month name according to the current locale.

%c The preferred date and time representation for the current locale.

%C  The century number (year/100) as a 2-digit integer.

%d The day of the month as a decimal number (range 01 to 31).

%D  Equivalent to ‘%m/%d/%y’.

%e Like ‘%d’, the day of the month as a decimal number, but a leading zero is replaced by a space.

%E Modifier: use alternative format, see below (see [conversion specs], page 63).

%F Equivalent to ‘%Y-%m-%d’ (the ISO 8601 date format).

%G The ISO 8601 year with century as a decimal number. The 4-digit year corresponding to the ISO week number (see ‘%V’). This has the same format and value as ‘%y’, except that if the ISO week number belongs to the previous or next year, that year is used instead.

%g Like ‘%G’, but without century, i.e., with a 2-digit year (00-99).

%h  Equivalent to ‘%b’.

%H The hour as a decimal number using a 24-hour clock (range 00 to 23).

%I The hour as a decimal number using a 12-hour clock (range 01 to 12).

%j The day of the year as a decimal number (range 001 to 366).

%k The hour (24-hour clock) as a decimal number (range 0 to 23); single digits are preceded by a blank. (See also ‘%H’.)

%l The hour (12-hour clock) as a decimal number (range 1 to 12); single digits are preceded by a blank. (See also ‘%I’.)

%m  The month as a decimal number (range 01 to 12).
%M  The minute as a decimal number (range 00 to 59).
%n  A newline character.

%O Modifier: use alternative format, see below (see [conversion specs], page 63).

%p Either ‘AM’ or ‘PM’ according to the given time value, or the corresponding strings for the current locale. Noon is treated as ‘pm’ and midnight as ‘am’.

%P Like ‘%p’ but in lowercase: ‘am’ or ‘pm’ or a corresponding string for the current locale.

%r The time in ‘a.m.’ or ‘p.m.’ notation. In the POSIX locale this is equivalent to ‘%I:%M:%S %p’.

%R The time in 24-hour notation (‘%H:%M’). For a version including the seconds, see ‘%T’ below.

%s The number of seconds since the Epoch, i.e., since 1970-01-01 00:00:00 UTC.

%S  The second as a decimal number (range 00 to 61).
%t  A tab character.
%T  The time in 24-hour notation (‘%H:%M:%S’).

%u The day of the week as a decimal, range 1 to 7, Monday being 1. See also ‘%w’.

%U The week number of the current year as a decimal number, range 00 to 53, starting with the first Sunday as the first day of week 01. See also ‘%V’ and ‘%W’.

%V The ISO 8601:1988 week number of the current year as a decimal number, range 01 to 53, where week 1 is the first week that has at least 4 days in the current year, and with Monday as the first day of the week. See also ‘%U’ and ‘%W’.

%w The day of the week as a decimal, range 0 to 6, Sunday being 0. See also ‘%u’.

%W The week number of the current year as a decimal number, range 00 to 53, starting with the first Monday as the first day of week 01.

%x The preferred date representation for the current locale without the time.

%X The preferred time representation for the current locale without the date.

%y The year as a decimal number without a century (range 00 to 99).

%Y The year as a decimal number including the century.

%z The time-zone as hour offset from GMT. Required to emit RFC822-conformant dates (using ‘%a, %d

%b %Y %H:%M:%S %z’)

%Z  The time zone or name or abbreviation.
%+  The date and time in date(1) format.
%%  A literal ‘%’ character.

Some conversion specifiers can be modified by preceding them by the ‘E’ or ‘O’ modifier to indicate that an alternative format should be used. If the alternative format or specification does not exist for the current locale, the behaviour will be as if the unmodified conversion specification were used. The Single Unix Specification mentions ‘%Ec’, ‘%EC’, ‘%Ex’, ‘%EX’, ‘%Ry’,

‘%EY’, ‘%Od’, ‘%Oe’, ‘%OH’, ‘%OI’, ‘%Om’, ‘%OM’, ‘%OS’, ‘%Ou’, ‘%OU’, ‘%OV’, ‘%Ow’,

‘%OW’, ‘%Oy’, where the effect of the ‘O’ modifier is to use alternative numeric symbols (say, roman numerals), and that of the ‘E’ modifier is to use a locale-dependent alternative representation.

Appendix B GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct

from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified

Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adja-

cent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice

giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections

and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License.

I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included in the Modified Version.

N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title with any Invariant Section.

O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties— for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document.

If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See

http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

       Copyright (C)  year  your name.
       Permission is granted to copy, distribute and/or modify this document
       under the terms of the GNU Free Documentation License, Version 1.3
       or any later version published by the Free Software Foundation;
       with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
       Texts.  A copy of the license is included in the section entitled ``GNU
       Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

         with the Invariant Sections being list their titles, with
         the Front-Cover Texts being list, and with the Back-Cover Texts
         being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index

This is a general index of all issues discussed in this manual.

$

$#. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

–

--count, rushlast . . . . . . . . . . . . . . . . . . . . . 55 --debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --debug, rush . . . . . . . . . . . . . . . . . . . . . . . . . 49 --dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 49 --file, rushlast . . . . . . . . . . . . . . . . . . . . . . 55 --file, rushwho . . . . . . . . . . . . . . . . . . . . . . 51 --format, rushlast . . . . . . . . . . . . . . . . . . . . 55 --format, rushwho. . . . . . . . . . . . . . . . . . . . 51 --forward, rushlast . . . . . . . . . . . . . . . . . . . 55 --help, rush . . . . . . . . . . . . . . . . . . . . . . . . . . 49 --help, rushlast . . . . . . . . . . . . . . . . . . . . . . 56 --help, rushwho . . . . . . . . . . . . . . . . . . . . . . 52 --lint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --lint, rush . . . . . . . . . . . . . . . . . . . . . . . . . . 49 --no-header, rushlast. . . . . . . . . . . . . . . . . 56 --no-header, rushwho . . . . . . . . . . . . . . . . 51 --security-check, rush . . . . . . . . . . . . . . 49 --show-default . . . . . . . . . . . . . . . . . . . . . . . 39 --show-default, rush . . . . . . . . . . . . . . . . . 49 --test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --test, rush . . . . . . . . . . . . . . . . . . . . . . . . . . 49 --trace, rush . . . . . . . . . . . . . . . . . . . . . . . . . 49 --usage, rush . . . . . . . . . . . . . . . . . . . . . . . . . 50 --usage, rushlast . . . . . . . . . . . . . . . . . . . . . 56 --usage, rushwho . . . . . . . . . . . . . . . . . . . . . 52 --user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --user, rush . . . . . . . . . . . . . . . . . . . . . . . . . . 49 --version, rush . . . . . . . . . . . . . . . . . . . . . . 49 --version, rushlast . . . . . . . . . . . . . . . . . . . 56 --version, rushwho. . . . . . . . . . . . . . . . . . . 52 -b, file system test . . . . . . . . . . . . . . . . . . . . . 19 -c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 -c, file system test . . . . . . . . . . . . . . . . . . . . . 19 -c, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -C, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -d. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 -d, file system test . . . . . . . . . . . . . . . . . . . . . 20 -d, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48, 49 -e, file system test . . . . . . . . . . . . . . . . . . . . . 20 -f, file system test . . . . . . . . . . . . . . . . . . . . . 20 -f, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -f, rushwho. . . . . . . . . . . . . . . . . . . . . . . . . . . 51 -F, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -F, rushwho. . . . . . . . . . . . . . . . . . . . . . . . . . . 51 -g, file system test . . . . . . . . . . . . . . . . . . . . . 20 -G, file system test . . . . . . . . . . . . . . . . . . . . 20 -h, file system test . . . . . . . . . . . . . . . . . . . . . 20 -h, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -h, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -h, rushwho. . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -H, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -H, rushwho. . . . . . . . . . . . . . . . . . . . . . . . . . . 51 -i, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -k, file system test . . . . . . . . . . . . . . . . . . . . . 20 -L, file system test . . . . . . . . . . . . . . . . . . . . . 20 -n, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -O, file system test . . . . . . . . . . . . . . . . . . . . 20 -p, file system test . . . . . . . . . . . . . . . . . . . . . 20 -r, file system test . . . . . . . . . . . . . . . . . . . . . 20 -s, file system test . . . . . . . . . . . . . . . . . . . . . 20 -S, file system test . . . . . . . . . . . . . . . . . . . . . 20 -t, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -T, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -u. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 -u, file system test . . . . . . . . . . . . . . . . . . . . . 20 -u, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -v, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -v, rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -v, rushwho. . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -w, file system test . . . . . . . . . . . . . . . . . . . . 20 -x, file system test . . . . . . . . . . . . . . . . . . . . . 20 -x, rush. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

A

accounting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 accounting database . . . . . . . . . . . . . . . . . . . 57 acct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 acct, dump attribute . . . . . . . . . . . . . . . . . 48 acct-dir-mode . . . . . . . . . . . . . . . . . . . . 14, 57 acct-file-mode . . . . . . . . . . . . . . . . . . . 14, 57 acct-umask . . . . . . . . . . . . . . . . . . . . . . . . 14, 57 actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 actions, system . . . . . . . . . . . . . . . . . . . . . . . . 27 all, dump attribute . . . . . . . . . . . . . . . . . . 48 all, include security flag. . . . . . . . . . . . . . . 13 argv, dump attribute . . . . . . . . . . . . . . . . . 48

76 GNU Rush – a restricted user shell

B

backreference interpretation. . . . . . . . . . . . . 8 backslash interpretation . . . . . . . . . . . . . . . . 8 basic regular expressions . . . . . . . . . . . . . . . 12

C

chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 checking file ownership . . . . . . . . . . . . . . . . 19 checking file type . . . . . . . . . . . . . . . . . . . . . . 19 chroot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

chroot_dir, dump attribute . . . . . . . . . . 48

clrenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

cmdline, dump attribute . . . . . . . . . . . . . 48

command . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 54 comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 conditions . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 18 config-error . . . . . . . . . . . . . . . . . . . . . . . . . 12 configuration file syntax . . . . . . . . . . . . . . . 10 configuration file, testing . . . . . . . . . . . . . . 47 cvs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

D

debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 debugging levels . . . . . . . . . . . . . . . . . . . . . . . 11 delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 dir_iwgrp, include security flag. . . . . . . . 13 dir_iwoth, include security flag. . . . . . . . 13 domain, localization . . . . . . . . . . . . . . . . . . . 34 dump mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 duration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

E

environ, dump attribute . . . . . . . . . . . . . 48

Environment . . . . . . . . . . . . . . . . . . . . . . . . . . 25 error messages. . . . . . . . . . . . . . . . . . . . . . . . . 12 evalenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 exit rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 expand-undefined. . . . . . . . . . . . . . . . . . . . . 11 expansion of undefined variables . . . . . . . 17 extended regular expressions . . . . . . . . . . . 12

F

fall-through . . . . . . . . . . . . . . . . . . . . . . . . . 29 fall-through rule . . . . . . . . . . . . . . . . . . . . . . . . 3 fall-through statement . . . . . . . . . . . . . . . . . 29 fallthrough . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 file ownership, checking . . . . . . . . . . . . . . . . 19 file type, checking . . . . . . . . . . . . . . . . . . . . . 19 fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 fork, dump attribute . . . . . . . . . . . . . . . . . 48 forked mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

G

g, transform flag . . . . . . . . . . . . . . . . . . . . . . 22 gecos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gid, dump attribute . . . . . . . . . . . . . . . . . . 48 git . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 git-receive-pack . . . . . . . . . . . . . . . . . . . . . . . . 44 git-shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 git-upload-pack . . . . . . . . . . . . . . . . . . . . . . . . 44 global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 19

groupwritabledir, include

security flag . . . . . . . . . . . . . . . . . . . . . . . . . 13

groupwritablefile, include

security flag . . . . . . . . . . . . . . . . . . . . . . . . . 13

H

home . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

home_dir, dump attribute . . . . . . . . . . . . 48

I

i, transform flag . . . . . . . . . . . . . . . . . . . . . . 22 i18n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 identifiers, configuration . . . . . . . . . . . . . . . . 8 `in', operator . . . . . . . . . . . . . . . . . . . . . . . 19 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 include-security. . . . . . . . . . . . . . . . . . . . . 13 indexing, words in command line . . . . . . 15 insert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 interactive . . . . . . . . . . . . . . . . . . . . . . . . . . 33 interactive access . . . . . . . . . . . . . . . . . . . . . . 33

interactive, dump attribute . . . . . . . . . 48

internationalization . . . . . . . . . . . . . . . . . . . . 33 iwgrp, include security flag . . . . . . . . . . . . 13 iwoth, include security flag . . . . . . . . . . . . 13

Concept Index 77

K

keepenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

L

l10n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 legacy syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 limiting number of

simultaneous sessions . . . . . . . . . . . . . . . . 29 limits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 link, include security flag . . . . . . . . . . . . . 13 locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 locale directory . . . . . . . . . . . . . . . . . . . . . . . . 34 locale name . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

locale, dump attribute . . . . . . . . . . . . . . 48

locale-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

localedir, dump attribute . . . . . . . . . . . 48

localization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 localization directives . . . . . . . . . . . . . . . . . . 35

M

map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 matching conditions . . . . . . . . . . . . . . . . . . . 18 message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 msgfmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

N

newgroup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 newgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 nologin-error . . . . . . . . . . . . . . . . . . . . . . . . . . 12

O

options, command line . . . . . . . . . . . . . . . . . 49 output formats . . . . . . . . . . . . . . . . . . . . . . . . 52 owner, include security flag . . . . . . . . . . . . 13

P

pid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 post-socket . . . . . . . . . . . . . . . . . . . . . . . . . . 31 prog, dump attribute . . . . . . . . . . . . . . . . . 48 program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

pw_dir, dump attribute . . . . . . . . . . . . . . 48
pw_gid, dump attribute . . . . . . . . . . . . . . 48
pw_name, dump attribute . . . . . . . . . . . . . 48
pw_uid, dump attribute . . . . . . . . . . . . . . 48

Q

quoted strings . . . . . . . . . . . . . . . . . . . . . . . . . . 8

R

regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 regular expressions . . . . . . . . . . . . . . . . . . . . 12 remopt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 request . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 15 rsync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 14, 54 rule tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 rule, fall-through . . . . . . . . . . . . . . . . . . . . . . . 3 rush, statement . . . . . . . . . . . . . . . . . . . . . . . 10 rush-po. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 rush.rc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 rushlast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 RUSHLAST_FORMAT . . . . . . . . . . . . . . . . . . . . . . 55 rushwho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 rushwho, command line options . . . . . . . . 51 RUSHWHO_FORMAT . . . . . . . . . . . . . . . . . . . . . . . 51

S

s-expression . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 scp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 22 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 sftp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 simultaneous sessions . . . . . . . . . . . . . . . . . . 29 sleep-time . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 start-time . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 stop-time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 svn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 syntax version statement . . . . . . . . . . . . . . 10 syntax, configuration files . . . . . . . . . . . . . . 10 syntax, legacy . . . . . . . . . . . . . . . . . . . . . . . . . . 7 system actions. . . . . . . . . . . . . . . . . . . . . . . . . 27 system-error . . . . . . . . . . . . . . . . . . . . . . . . . 12

78 GNU Rush – a restricted user shell

T

tab. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 tag, rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 tcpmux. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 test mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 testing configuration file . . . . . . . . . . . . . . . 47 text-domain . . . . . . . . . . . . . . . . . . . . . . . . . . 35

text_domain, dump attribute . . . . . . . . . 48

textual domain . . . . . . . . . . . . . . . . . . . . . . . . 34 tilde expansion. . . . . . . . . . . . . . . . . . . . . 28, 36 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 time formats, for

--time-format option . . . . . . . . . . . . . . . 61 trap rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

U

uid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

umask, dump attribute. . . . . . . . . . . . . . . . 48

undefined variable, expansion . . . . . . . . . . 17 unquoted strings. . . . . . . . . . . . . . . . . . . . . . . . 8 unset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 usage-error . . . . . . . . . . . . . . . . . . . . . . . . . . 12 user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 53 utmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 utmp file, accounting database . . . . . . . . . 57

V

vars, dump attribute . . . . . . . . . . . . . . . . . 48

W

word splitting . . . . . . . . . . . . . . . . . . . . . . . . . 15

                                        worldwritabledir, include

security flag . . . . . . . . . . . . . . . . . . . . . . . . . 13

                                        worlwritablefile, include

security flag . . . . . . . . . . . . . . . . . . . . . . . . . 13 wtmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 wtmp file, accounting database . . . . . . . . . 57

X

x, transform flag . . . . . . . . . . . . . . . . . . . . . . 22